OpenStudy (anonymous):

Let \(\Large f(x) = log(1-x)\) b) Expand \(\Large f^{(k)}(x)\) in a power series around \(\Large x=0\) and determine its radius of convergence.

OpenStudy (helder_edwin):

the superscript k is the order of derivative or the number of times the function is composed \[ f^k(x)\neq f^{(k)}(x) \] you should be careful with notation

OpenStudy (anonymous):

i have edited thx ;)

OpenStudy (helder_edwin):

ok. so \[ \LARGE f'(x)=\frac{-1}{1-x} \]

OpenStudy (helder_edwin):

then \[ \LARGE f''(x)=\frac{-(-1)(-1)}{(1-x)^2}=\frac{-1}{(1-x)^2} \] agree?

OpenStudy (anonymous):

hmm it looks good..

OpenStudy (anonymous):

i have question but i want to ask when you finish..

OpenStudy (helder_edwin):

then \[ \LARGE f'''(x)=\frac{-(-1)2(1-x)(-1)}{(1-x)^4}=\frac{-2}{(1-x)^3} \] \[ \LARGE f^{(4)}(x)=\frac{-(-2)3(1-x)^2(-1)}{(1-x)^6}=\frac{-6}{(1-x)^4} \]

OpenStudy (helder_edwin):

i think you can infer the general formula for the k-th derivative from all this

OpenStudy (anonymous):

which general formula you mean, can you post a link ?

OpenStudy (helder_edwin):

i mean \[ \LARGE f^{(k)}(x)=\frac{-(k-1)!}{(1-x)^k} \] for k>=1